Hel Given the following information: Percent of capital structure: Preferred #toek Comnon equity (retained earnings) Debt 358 45 20 eBook Adltionalinformation: Corporate tax rate Dividend, preferred Dividend, expected common $ 3.50 Price, preferzed Growth rate Bond yield rlotation coat, preterred 4.50 Price, common $ 8.00 $110.00 6 85.00 Calculate the weighted average cost of capital for Digital Processing Inc. (Do not round intermediate ca answers as a percent rounded to 2 decimal places) intermediate calculations. Input your Weighted Cost Debt Preferred stock Common equity (retained earnings) Weighted average cost of capital く Prey 5015 Ⅲ Next >
Answer #1
A B C D E F G H I J
2
3 Formula for WACC is given as:
4 WACC = r(E) × w(E) + r(P) × w(P)+r(D) × (1 – t) × w(D)
5 Where, r(E), r(P) and r(D) are cost of equity, preferred stock and cost of debt, w(E), w(P) and W(D)
6 are weight of equity, preferred stock and debt and t is the tax rate
7 Calculation of cost of debt:
8 Cost of debt will be the yield to maturity of the bond can be calculated as follows:
9
10 Yield to maturity of the bond 6.00%
11
12 Hence Cost of debt is 6.00%
13
14 Calculation of cost of common equity using dividend growth:
15
16 As per dividend growth model, Price of share is the present value of all future dividends discounted at cost of equity.
17 Current Price of Share (P) =Div1)/(r-g)
18 Where Div1 is the dividend paid next year, r is cost of equity and g is perpetual growth rate of dividend.
19
20 Cost of equity can be calculated as below:
21 Cost of equity, r(E) = (Div1/P)+g
22
23 Given the following data:
24 Dividend (Div1) $3.50
25 Price (P) $85
26 Growth rate (g) 6.0%
27
28 From Dividend growth model,
29 Cost of equity, r(E) = (Div1/P)+g
30 Cost of equity= 10.12% =(D24/D25)+D26
31
32 Hence cost of using dividend growth model is 10.12%
33
34 Calculation of Cost of preferred stock and market value of preferred stock:
35
36 Annual Dividend of preferred stock $8.00
37 Current Price $110.00
38 Floatation cost (F) $4.50
39 Cost of preferred stock =Preferred Dividend/(Current Price - Preferred Stock)
40 7.58% =D36/(D37-D38)
41
42 Hence cost of Preferred Stock is 7.58%
43
44 Calculation of Weighted Cost
45
46 Tax Rate 30%
47
48 Source of capital Capital Structure Cost Weighted Cost
49 Debt 20.00% 6.00% 0.84% =E49*(1-D46)*D49
50 Common Stock 45.00% 10.12% 4.55% =E50*D50
51 Preferred Stock 35.00% 7.58% 2.65% =E51*D51
52 WACC 8.05% =SUM(F49:F51)
53